OpenStudy (anonymous):

I GIVE MEDALS!!!!!!!!! I just need to know how to do this. Find each product. 5/6*10

OpenStudy (anonymous):

\[\frac{ 5 }{ 6 }\]\[\times 10\]

OpenStudy (anonymous):

if you have ten lots of it you will have: \[\frac{ 5 \times 10 }{ 6 }=?\]

OpenStudy (anonymous):

I'm confused.

OpenStudy (anonymous):

by what?

OpenStudy (anonymous):

\[\frac{ 50 }{ 6 }\]?

OpenStudy (anonymous):

Okay well you can cancel it down by finding a common factor between the top and the bottom, so what divides into both 6 and 50?

OpenStudy (anonymous):

I got a weird answer i got 25/3

OpenStudy (anonymous):

^^ that's right

OpenStudy (anonymous):

it is yes

OpenStudy (anonymous):

(but I was trying to get them to use the method so they understand it)

OpenStudy (anonymous):

So I multiply the numerator by the other number and then what?

OpenStudy (anonymous):

I got it right

OpenStudy (anonymous):

then you just cancel down to get the fraction in it's simplest terms

OpenStudy (anonymous):

What do you mean by cancel down?

OpenStudy (anonymous):

actually what sara said

OpenStudy (anonymous):

you find a common factor between the numerator and denominator, for example both 6 and 50 are divisible by 2

OpenStudy (anonymous):

divide the numbers and multiply it by 10

OpenStudy (anonymous):

Oh okay! Thanks! If I could give you both I would!

OpenStudy (anonymous):

Medals

OpenStudy (anonymous):

don't worry about it, i'm not fussed.

OpenStudy (anonymous):

i would give it to sara

OpenStudy (anonymous):

as long as you understand that's fine

OpenStudy (anonymous):

Thanks you two!

OpenStudy (anonymous):

she explained it more

OpenStudy (anonymous):

Still you helped. Thanks!

OpenStudy (anonymous):

your welcome, if you have anymore questions just ask ok

OpenStudy (anonymous):

Okay thanks!